condense 5 log 5 + 5 log 11 + 20 log 6

Condense each expression to a single 1 I I 5) 5log 5 + 5log 11 + 20 log 6
Answer #1

5 log 5 + 5log 11 + 20 loge = log (55) + log (115) +log (6²) [ log x=nlogx] log (5²x11x6 620) [ log(x) = log(x) +10960 Log (